OpenStudy (anonymous):

PLEASE HELP !!!!!!!! Write an algebraic expression for each phrase. 11 decreased by a variable n. a variable x minus 4. the sum of a variable x and 7.

OpenStudy (cwrw238):

11 - n x - 4 x + 7

OpenStudy (anonymous):

11-n x-4 x+7

OpenStudy (anonymous):

THANK YOU SO MUCH

OpenStudy (anonymous):

Now help me in a few more, please.

OpenStudy (anonymous):

25 divides by a variable n

OpenStudy (cwrw238):

25/n

OpenStudy (anonymous):

Now off to a different topic in math

OpenStudy (anonymous):

Estimate each expression . Use the order of operations when necessary.

OpenStudy (anonymous):

a + 1, for = 10

OpenStudy (cwrw238):

10 + 1 = 11 (for a = 10)

OpenStudy (anonymous):

okay , b - 3 , for b = 3

OpenStudy (cwrw238):

3-3 = 0

OpenStudy (anonymous):

9 + c -5 , for c = 7

OpenStudy (anonymous):

PLEASE HELP MelindaR

OpenStudy (anonymous):

9 + c -5 , for c = 7

OpenStudy (anonymous):

9 + 7 - 5 = ?

OpenStudy (anonymous):

what's 9 + 7?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

and 16 - 5?

OpenStudy (anonymous):

11

OpenStudy (anonymous):

there you go

OpenStudy (anonymous):

So c equals to 11

OpenStudy (anonymous):

guide me through the next one , okay

OpenStudy (anonymous):

d * 8 , for d = 9

OpenStudy (anonymous):

9 * 8 = ?

OpenStudy (anonymous):

72

OpenStudy (anonymous):

good

OpenStudy (anonymous):

so the answer is 72?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

( e + 4 ) / 2 , for e = 8

OpenStudy (anonymous):

( 8 + 4) / 2 = ?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

yeah that's it

OpenStudy (anonymous):

but, wait the question says e = 8

OpenStudy (anonymous):

whats 6

OpenStudy (anonymous):

( 6 + 4 ) / 2 , for e = 6 is that right?

OpenStudy (anonymous):

( 8 + 4) / 2 = 6 because 8 + 4 is equal 12. and 12/2 is equal 6. got it?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

okay , last question

OpenStudy (anonymous):

5 + 8 * f , for f = 2

OpenStudy (anonymous):

replace f for 2, can you do that?

OpenStudy (anonymous):

i don't know

OpenStudy (anonymous):

type the equation and instead of f you put 2

OpenStudy (anonymous):

so , the expression would look like this 5 + 8 * 2 , for f = 2

OpenStudy (anonymous):

yes, very good

OpenStudy (anonymous):

now solve

OpenStudy (anonymous):

first 8 * 2

OpenStudy (anonymous):

16

OpenStudy (anonymous):

then + 5 is 21

OpenStudy (anonymous):

yeah correct

OpenStudy (anonymous):

so , now the equation looks like ?

OpenStudy (anonymous):

5 + 8 * f , for f = 2 5 + 8 * 2 5 + 16 21

OpenStudy (anonymous):

thanks a lot

OpenStudy (anonymous):

anytime

OpenStudy (anonymous):

okay